OpenStudy (anonymous):

If two balls are drawn without replacement from a box containing 3 red balls, 2 black balls and one green ball, what is the probability of getting 2 red balls? Can i have some help with this please?????

OpenStudy (anonymous):

there are 3+2+1=6 balls in the box if you pick two ball at once it should be 2/6=1/3 if you pick one ball at once and pick twice it should be (3/6)*(2/5)=1/5